OpenStudy (anonymous):

In a circle of radius 60 inches, a central angle of 35 will intersect the circle forming an arc of length ______

OpenStudy (caozeyuan):

35 degrees right?

OpenStudy (anonymous):

yes

OpenStudy (caozeyuan):

so what is 35 degrees in radians

OpenStudy (nikato):

or this is how i think of it. first find circumference. remember C=pi(d) and you know d=diameter=120 so circumference is 120pi

OpenStudy (anonymous):

Idon'tknow

OpenStudy (caozeyuan):

so 35/360=X/2pi, can you solve for x?

OpenStudy (nikato):

and 360 degrees in a circle and 35/360 = 7/72 so now whats 7/72 (120pi)

OpenStudy (anonymous):

36.65 inches